OpenStudy (anonymous):

Media service charges % 20 for a phone and $30/moth for its economy plan. The equation c-30t+20 describe c= 30t+20 describes the total cost, of the operating a media service phone for t months. The total cost for 7 months of service $____

OpenStudy (saifoo.khan):

i doubt that is $20.

OpenStudy (anonymous):

i meant to say $ 20 for a phone

OpenStudy (anonymous):

30/month

jimthompson5910 (jim_thompson5910):

c= 30t+20 c= 30(7)+20 c= 210+20 c= 230 The total cost for 7 months of service is $230